# Dissecting Lepton Number Violation in the Left-Right Symmetric Model: $0\nu\beta\beta$ decay, Møller scattering, and collider searches

### v1

@@ V Results and Discussion
Notice that these parameters completely determine the right-handed neutrino mass $M_{N_{R}}$ by combining Eqs. (19) and (46). Since parity-violating Møller scattering is only sensitive to the subset of parameters $\{M_{\delta_{R}^{\pm\pm}},(f_{R})_{ee}\}$ (cf. Eq. (45) and subsequent discussion), we choose different benchmark scenarios for $M_{W_{R}}$ and the left-right mixing angle $\zeta$. These values were selected to be consistent with the constraints discussed in Sec. IV.
@@
Using the expressions in Sec. IV.1, we obtain constraints from the ${0\nu\beta\beta}$-decay searches in Figs. 3 and 4 for the active neutrino masses in the normal hierarchy (NH)  [^4] There are no significant differences when the inverted hierarchy is chosen, so we omit the results for the sake of brevity.. A parameter scan over $M_{\delta_{R}^{\pm\pm}}$ and $|(f_{R})_{ee}|$ is performed to predict $T_{1/2}^{0\nu}$ and compare it with the most recent KamLAND-Zen results ($T_{1/2}^{0\nu}>2.3\times 10^{26}$ years at 90% C.L. Abe et al. (2022)) to obtain the excluded parameter regions shown as red points. The solid red lines bound these regions, and they correspond to theoretical estimates of the ${0\nu\beta\beta}$ amplitude based on the different contributions (doubly-charged scalar, active neutrino, and sterile neutrino) in Eq. (25), as we will explain in the next paragraphs. We use the same theoretical estimate to draw the exclusion boundary using the ton-scale prospects that are expected to be sensitive to the half-life of $10^{28}$ years Agostini et al. (2017); these results are shown as dashed magenta lines.
@@
Fig. 3 shows two separate regions that are consistent with the observed ${0\nu\beta\beta}$-decay constraints. The top region accommodates large Yukawa couplings and heavy doubly-charged scalars, while the lower region is independent of the doubly-charged scalar mass and is consistent with small couplings. The gap between these two regions corresponds to the transition between a heavy and a light right-handed neutrino description in the amplitude in Eq. (IV.1). To better understand these results, we will analyze the different contributions to the amplitude and identify the ranges where the individual contributions shown in Fig. 1 dominate.
@@
Figure 3: Parameter scan of ${0\nu\beta\beta}$-decay constraints on the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and the Yukawa coupling $|(f_{R})_{ee}|$ in the type-II seesaw scenario of the LRSM with zero (left) non-zero (right) $W_{L}-W_{R}$ mixing. The active neutrino masses are assumed in the normal hierarchy (NH), and the right-handed $W$ boson mass is chosen as $M_{W_{R}}=7\,\mathrm{TeV}$. We take $\xi=0$ and $0.35$ for the left- and right-handed plots, respectively, and use Eq. (21) to characterize the results in terms of the physical mixing angle $\zeta$. The right-handed neutrino mass $M_{N_{R}}$ is also displayed as a secondary vertical axis. The red points (regions above or between the solid red lines) are excluded by the most recent KamLAND-Zen results Abe et al. (2022). The solid red lines correspond to the exclusion limits calculated using an analytical estimation (see text for discussion). The dashed magenta lines correspond to the same exclusion limits using the prospect by ton-scale experiments Agostini et al. (2017).
Figure 4: The figure shows six panels, each of them similar to Fig. 3, emphasizing the effects of varying the $M_{W_{R}}$ mass. The three benchmarks displayed from left to right correspond to $M_{W_{R}}=7\,\mathrm{TeV}$, $15\,\mathrm{TeV}$ and $25\,\mathrm{TeV}$. The top row depicts the scenario with zero $W_{L}-W_{R}$ mixing, while the bottom row shows the case where a non-zero $W_{L}-W_{R}$ mixing was assumed.
@@
The specific dependence of $T_{1/2}^{0\nu}$ as a function of $|(f_{R})_{ee}|$ can be determined by analyzing the different contributions to the amplitude in Eq. (IV.1). There are four regimes to be considered:
@@
  • Doubly-charged scalar regime. For a large Yukawa coupling, the doubly-charged scalar contribution to the amplitude dominates the inverse half-life in Eq. (25). By analyzing diagrams (e) and (f) in Fig. 1, we notice that the inverse half-life resulting after integrating out $W_{R}$ and $\delta_{R}^{\pm\pm}$ will be proportional to
@@
    where the second relation in Eq. (19) has been used. The overall factor is given by the product of the phase-space factor, NMEs, and low-energy constants. It also depends on the $W_{L}-W_{R}$ mixing and indicates that in this regime, for a larger $M_{\delta_{R}^{\pm\pm}}$, a larger $|(f_{R})_{ee}|$ is allowed by the constraints of ${0\nu\beta\beta}$-decay experiments, shown as the diagonal boundaries in Figs. 3 and 4. It is worth noting that this region of the parameter space aligns with the findings in Ref. Dev et al. (2018), which focused only on the effects of the doubly-charged scalar as new physics contributions to ${0\nu\beta\beta}$-decay. Additionally, when the left-right mixing is turned on, the contribution to the inverse half-life increases so that the region is shrunk to the right.
@@
    where we have used the relation in Eq. (46) to obtain the extra power of $M_{W_{R}}$ in the amplitude’s denominator. This contribution is independent of $M_{\delta_{R}^{\pm\pm}}$ and denoted by the upper horizontal boundary in Figs. 3 and 4. The contribution to the inverse half-life also increases if the left-right mixing is turned on so that a larger $|(f_{R})_{ee}|$ is preferred by comparing the left and right panels in Figs. 3 and 4.
@@
Figure 5: Inverse half-life $(T^{0\nu}_{1/2})^{-1}$ as a function of the Yukawa coupling $|(f_{R})_{ee}|$ for zero (left) and non-zero $W_{L}-W_{R}$ mixing (right). We fix $\xi=0$ and $\xi=0.35$ for the left and right plots, respectively. We also fix the mass of the doubly-charged scalar at $M_{\delta_{R}^{\pm\pm}}=2000\,\mathrm{GeV}$. The three curves in the figure show the results for $M_{W_{R}}=7\,\mathrm{TeV}$ (blue), $M_{W_{R}}=15\,\mathrm{TeV}$ (yellow), and $M_{W_{R}}=25\,\mathrm{TeV}$ (green). We also show the corresponding values of the mixing angle $\zeta$. The inverse half-life considering only the light left-handed neutrino contributions is shown in red. The latest KamLAND-Zen exclusion limits are represented by the grey shaded area, with the future ton-scale experiment bounds indicated for comparison.
@@
Figure 6: Combined sensitivities to the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and the Yukawa coupling $|(f_{R})_{ee}|$ in the ${0\nu\beta\beta}$-decay, MOLLER, and collider experiments in the type-II seesaw scenario of the LRSM. The normal hierarchy is assumed. We fix the mass of the $W_{R}$ boson to $M_{W_{R}}=7\,\mathrm{TeV}$ and present scenarios with zero (left) and non-zero (right) $W_{L}-W_{R}$ mixing. The right-handed neutrino mass $M_{N_{R}}$ is also displayed as a secondary vertical axis. The MOLLER prospect Benesch et al. (2014) is shown as a dashed blue line, and ${0\nu\beta\beta}$ decay limits from current and future experiments are shown as solid red and dashed magenta lines, respectively. The figure also displays the direct searches at the LHC (solid gray), pair production prospect at the FCC-hh (dashed gray), and Bhabha scattering limits at LEP (solid orange), the CEPC (dashed green), and the FCC-ee (dashed dark green) prospects. The ${0\nu\beta\beta}$-decay constraint by the KamLAND-Zen experiment from Ref. Dev et al. (2018) (dash-dotted black) is also included for comparison.
@@
To understand the gap in Fig. 3, let us consider $(T_{1/2}^{0\nu})^{-1}$ as a function of $|(f_{R})_{ee}|$ for $M_{\delta_{R}^{\pm\pm}}=2\,\mathrm{TeV}$, $M_{W_{R}}=7\,\mathrm{TeV}$, and zero/non-zero $W_{L}-W_{R}$ mixing, which corresponds to take vertical slices in both panels of Fig. 3.
@@
We present these results in Fig. 5 as a solid blue line. For the left and right panels, we consider $\xi=0$ and 0.35, respectively, and use Eq. (21) to characterize the results in terms of the physical mixing angle $\zeta$ and $M_{W_{R}}$ in case of a non-zero $W_{L}-W_{R}$ mixing. To show $(T_{1/2}^{0\nu})^{-1}$ in the standard mechanism, a red dashed curve is added. The grey shaded area represents the latest KamLAND-Zen exclusion limit, with the future ton-scale experiment bound indicated by the dashed black link for comparison.
@@
From Fig. 5, we find that in the large coupling region, where the amplitude is dominated by the doubly-charged scalar contribution, $(T_{1/2}^{0\nu})^{-1}$ decreases as $|(f_{R})_{ee}|$ is reduced, remaining consistent with current experimental bounds. We then observe a transition to a second regime, where the heavy right-handed neutrino contribution dominates the amplitude. In this regime, $(T_{1/2}^{0\nu})^{-1}$ increases as $|(f_{R})_{ee}|$ decreases. Consequently, for values of $|(f_{R})_{ee}|\lesssim 10^{-3}$, the model is excluded by current experimental data. However, around $|(f_{R})_{ee}|\sim 10^{-5}$, the right-handed neutrino becomes sufficiently light to alter the functional dependence of $(T_{1/2}^{0\nu})^{-1}$, causing it to decrease again as $|(f_{R})_{ee}|$ is further reduced. As expected, for values of $|(f_{R})_{ee}|\lesssim 10^{-7}$, $(T_{1/2}^{0\nu})^{-1}$ becomes compatible with current experimental bounds again.
[…]
@@
Figure 7: Same as in Fig. 6, combined constraints from LNC and ${0\nu\beta\beta}$-decay experiments on the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and the Yukawa coupling $|(f_{R})_{ee}|$ in the parity-violating LRSM with type-II seesaw dominance in the normal hierarchy with $M_{W_{R}}=15\,\mathrm{TeV}$ for zero (left) and non-zero (right) $W_{L}-W_{R}$ mixing.
@@
To illustrate the complementary sensitivities of ${0\nu\beta\beta}$-decay searches and LNC processes in both low-energy experiments and high-energy colliders, we present results based on current constraints and future prospects, as described in Sec. III and Sec. IV. The experimental sensitivities to $M_{\delta_{R}^{\pm\pm}}$ and $|(f_{R})_{ee}|$ are shown in Figs. 6 and 7 for zero and non-zero mixing angles. In each plot, the vertical axis on the right gives the values of $M_{N_{R}}$ that corresponds to $|(f_{R})_{ee}|$ using Eq. (46).
@@
In these two figures, the ${0\nu\beta\beta}$ decay exclusion limits are shown as red solid lines and correspond to the doubly-charged scalar (top diagonal), heavy right-handed neutrino (upper horizontal), and light right-handed neutrino (lower horizontal) contributions to the amplitude. They were obtained using the analytical expressions in Eqs. (47)–(49). The shaded red regions limited by those lines are excluded by the KamLAND-Zen experiment. Similar prospects for ton-scale scale experiments are included in dashed magenta lines. For comparison purposes, we also show the ${0\nu\beta\beta}$ decay exclusion limit derived in Ref. Dev et al. (2018) with the left-right mixing angle $\zeta=0$. Rescaling their limit to accommodate the latest KamLAND-Zen results and modifying the corresponding values of $M_{W_{R}}$ [^5] Ref. Dev et al. (2018) assumed a fixed value of $M_{W_{R}}=3.3\,\mathrm{TeV}$, which has been excluded by the LHC searches at Run-2 phase Aad et al. (2023a); Tumasyan et al. (2022a, b). yields the black dot-dashed curve in Fig. 6. Notice how the use of an EFT description provides a stronger constraint in the ${0\nu\beta\beta}$-decay sensitivity compared to the existing, state-of-the-art literature Dev et al. (2018).
@@
We also include collider searches in Figs. 6 and 7. The current indirect LEP exclusion limit for Bhabha scattering is shown as a solid orange line, with its corresponding exclusion region shaded in orange. The prospects for Bhabha scattering at the CEPC and the FCC-ee are shown as dashed dark green and light green lines, respectively, using the rescaling factor in Eq. (22). Notice that the difference in luminosities between both accelerators makes the CEPC bound stronger. These collider searches will reach into the “funnel” portion of the parameter space the ${0\nu\beta\beta}$ decay cannot access with current envisioned experiments. Direct searches for right-handed doubly-charged scalar $\delta_{R}^{\pm\pm}$ at hadron colliders are also shown for both current (LHC, grey solid line) and future (FCC-hh, grey dashed line) experiments, as discussed in Sec. III. They are sensitive to $M_{\delta_{R}^{\pm\pm}}\lesssim 5\,\mathrm{TeV}$ irrespective of the Yukawa coupling $|(f_{R})_{ee}|$. If we conceive a scenario where a positive event in ton-scale ${0\nu\beta\beta}$ decay experiments is seen but both future leptons and hadrons colliders do not get a corresponding signal, then we can pin down the underlying ${0\nu\beta\beta}$-decay mechanism to be dominated by the right-handed neutrino exchange and the corresponding parameter region would be a horizontal fringe of smaller $|(f_{R})_{ee}|$ in Figs. 6 and 7.
VI Conclusion

### v2

@@ 5 Results and Discussion
Notice that these parameters completely determine the right-handed neutrino mass $M_{N_{R}}$ by combining Eqs. (17) and (44), $M_{N_{R}}=(2/g_{R})\,M_{W_{R}}\,|(f_{R})_{ee}|$. We do not restrict ourselves to a specific underlying model but adopt the assumption $g_{R}=2g_{L}/3$ inspired by Ref. [Deppisch:2014zta] in the numerical analysis for illustrative purposes. Results for other scenarios, such as $g_{R}=g_{L}$, can be easily derived within our framework. Since parity-violating Møller scattering is only sensitive to the subset of parameters $\{M_{\delta_{R}^{\pm\pm}},|(f_{R})_{ee}|\}$ (cf. Eq. (43) and subsequent discussion), we choose different benchmark scenarios for $M_{W_{R}}$ and the left-right mixing angle $\zeta$. These values were selected to be consistent with the constraints discussed in Sec. 4.
@@
Using the expressions in Sec. 4.1, we obtain constraints from the $0\nu\beta\beta$}-decay searches in Fig. 3 for the active neutrino masses in the normal hierarchy (NH)  [^5] There are no significant differences when the inverted hierarchy is chosen, so we omit the results for the sake of brevity.. A parameter scan over $M_{\delta_{R}^{\pm\pm}}$ and $|(f_{R})_{ee}|$ is performed to predict $T_{1/2}^{0\nu}$ and compare it with the most recent KamLAND-Zen results ($T_{1/2}^{0\nu}>2.3\times 10^{26}$ years at 90% C.L. [KamLAND-Zen:2022tow]) to obtain the excluded parameter regions shown as red points. The solid red lines bound these regions, and they correspond to theoretical estimates of the $0\nu\beta\beta$} amplitude based on the different contributions (doubly-charged scalar, active neutrino, and sterile neutrino) in Eq. (23), as we will explain in the next paragraphs. We use the same theoretical estimate to draw the exclusion boundary, using ton-scale prospects that are expected to be sensitive to a half-life of $10^{28}$ years [Agostini:2017jim]; these results are shown as dashed magenta lines.
@@
The top row in Fig. 3 shows two separate regions that are consistent with the observed $0\nu\beta\beta$}-decay constraints. The top region accommodates large Yukawa couplings and heavy doubly-charged scalars, while the lower region is independent of the doubly-charged scalar mass and is consistent with small couplings. The gap between these two regions corresponds to the transition between a heavy and a light right-handed neutrino description in the amplitude in Eq. (4.1). To better understand these results, we will analyze the different contributions to the amplitude and identify the ranges where the individual contributions shown in Fig. 1 dominate.
@@
Figure 3: Parameter scan of $0\nu\beta\beta$}-decay constraints on the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and the Yukawa coupling $|(f_{R})_{ee}|$ in the LRSM with zero (left column) and non-zero (right column) $W_{L}-W_{R}$ mixing. The active neutrino masses are assumed in the normal hierarchy (NH), and the right-handed $W$ boson mass is chosen as $M_{W_{R}}=7\,\mathrm{TeV}$ (top row) and $M_{W_{R}}=15\,\mathrm{TeV}$ (bottom row). We take $\xi=0$ and $0.35$ for the left- and right-handed plots, respectively, and use Eq. (19) to characterize the results in terms of the physical mixing angle $\zeta$. The right-handed neutrino mass $M_{N_{R}}$ is also displayed as a secondary vertical axis. The red points (regions above or between the solid red lines) are excluded by the most recent KamLAND-Zen results [KamLAND-Zen:2022tow]. The solid red lines correspond to the exclusion limits calculated using an analytical estimation (see text for discussion). The dashed magenta lines correspond to the same exclusion limits using the prospect by ton-scale experiments [Agostini:2017jim].
@@
The specific dependence of $T_{1/2}^{0\nu}$ as a function of $|(f_{R})_{ee}|$ can be determined by analyzing the different contributions to the amplitude in Eq. (4.1). There are four regimes to be considered:
@@
  • Doubly-charged scalar regime. For a large Yukawa coupling, the doubly-charged scalar contribution to the amplitude dominates the inverse half-life in Eq. (23). By analyzing diagrams (e), (f), and (g) in Fig. 1, we notice that the inverse half-life resulting after integrating out $W_{R}$ and $\delta_{R}^{\pm\pm}$ will be proportional to
@@
    where the second relation in Eq. (17) has been used. The overall factor is given by the product of the phase-space factor, NMEs, and low-energy constants. It also depends on the $W_{L}-W_{R}$ mixing and indicates that in this regime, for a larger $M_{\delta_{R}^{\pm\pm}}$, a larger $|(f_{R})_{ee}|$ is allowed by the constraints of $0\nu\beta\beta$}-decay experiments, shown as the diagonal boundaries in Fig. 3. It is worth noting that this region of the parameter space aligns with the findings in Ref. [Dev:2018sel], which focused only on the effects of the doubly-charged scalar as new physics contributions to $0\nu\beta\beta$}-decay. Additionally, when the left-right mixing is turned on, the contribution to the inverse half-life increases so that the region is shrunk to the right.
@@
    where we have used the relation in Eq. (44) to obtain the extra power of $M_{W_{R}}$ in the amplitude’s denominator. This contribution is independent of $M_{\delta_{R}^{\pm\pm}}$ and denoted by the upper horizontal boundary in Fig. 3. The contribution to the inverse half-life also increases if the left-right mixing is turned on so that a larger $|(f_{R})_{ee}|$ is preferred by comparing the left and right panels in Fig. 3.
@@
Figure 4: Inverse half-life $(T^{0\nu}_{1/2})^{-1}$ as a function of the Yukawa coupling $|(f_{R})_{ee}|$ for zero (left) and non-zero $W_{L}-W_{R}$ mixing (right). We fix $\xi=0$ and $\xi=0.35$ for the left and right plots, respectively. We also fix the mass of the doubly-charged scalar at $M_{\delta_{R}^{\pm\pm}}=2000\,\mathrm{GeV}$. The two curves in the figure show the results for $M_{W_{R}}=7\,\mathrm{TeV}$ (blue) and $M_{W_{R}}=15\,\mathrm{TeV}$ (yellow). We also show the corresponding values of the mixing angle $\zeta$. The inverse half-life considering only the light left-handed neutrino contributions is shown in red. The latest KamLAND-Zen exclusion limits are represented by the grey shaded area, with the future ton-scale experiment bounds indicated for comparison.
@@
Figure 5: Combined sensitivities to the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and the Yukawa coupling $|(f_{R})_{ee}|$ in the $0\nu\beta\beta$}-decay, MOLLER, and collider experiments in the LRSM. The normal hierarchy is assumed. We fix the mass of the $W_{R}$ boson to $M_{W_{R}}=7\,\mathrm{TeV}$ and present scenarios with zero (left) and non-zero (right) $W_{L}-W_{R}$ mixing. The right-handed neutrino mass $M_{N_{R}}$ is also displayed as a secondary vertical axis. The MOLLER prospect [MOLLER:2014iki] is shown as a dashed blue line, and $0\nu\beta\beta$} decay limits from current and future experiments are shown as solid red and dashed magenta lines, respectively. The figure also displays the direct searches at the LHC (solid gray), pair production prospect at the FCC-$h$ (dashed gray), and Bhabha scattering limits at LEP (solid orange), the CEPC (dashed green), and the FCC-ee (dashed dark green) prospects.
@@
To understand the gap in the top row in Fig. 3, let us consider $(T_{1/2}^{0\nu})^{-1}$ as a function of $|(f_{R})_{ee}|$ for $M_{\delta_{R}^{\pm\pm}}=2\,\mathrm{TeV}$, $M_{W_{R}}=7\,\mathrm{TeV}$, and zero/non-zero $W_{L}-W_{R}$ mixing, which corresponds to taking vertical slices in both panels of the top row in Fig. 3.
@@
We present these results in Fig. 4 as a solid blue line. For the left and right panels, we consider $\xi=0$ and 0.35, respectively, and use Eq. (19) to characterize the results in terms of the physical mixing angle $\zeta$ and $M_{W_{R}}$ in the case of a non-zero $W_{L}-W_{R}$ mixing. To show $(T_{1/2}^{0\nu})^{-1}$ in the standard mechanism, a red dashed curve is added. The grey shaded area represents the latest KamLAND-Zen exclusion limit, with the future ton-scale experiment bound indicated by the dashed gray line for comparison.
@@
From Fig. 4, we find that in the large coupling region, where the amplitude is dominated by the doubly-charged scalar contribution, $(T_{1/2}^{0\nu})^{-1}$ decreases as $|(f_{R})_{ee}|$ is reduced, remaining consistent with current experimental bounds. We then observe a transition to a second regime, where the heavy right-handed neutrino contribution dominates the amplitude. In this regime, $(T_{1/2}^{0\nu})^{-1}$ increases as $|(f_{R})_{ee}|$ decreases. Consequently, for values of $|(f_{R})_{ee}|\lesssim 10^{-3}$, the model is excluded by current experimental data. However, around $|(f_{R})_{ee}|\sim 10^{-5}$, the right-handed neutrino becomes sufficiently light to alter the functional dependence of $(T_{1/2}^{0\nu})^{-1}$, causing it to decrease again as $|(f_{R})_{ee}|$ is further reduced. As expected, for values of $|(f_{R})_{ee}|\lesssim 10^{-7}$, $(T_{1/2}^{0\nu})^{-1}$ becomes compatible with current experimental bounds again  [^6] For such a small Yukawa coupling $(f_{R})_{ee}$, the mass of the right-handed neutrino that couples to the electron is $M_{N_{R}}\gtrsim\mathcal{O}({\rm MeV})$. One might be concerned about whether the constraints from various experimental searches [Bolton:2019pcu, Alves:2023znq] can be evaded. In the type-I seesaw mechanism, the active-sterile mixing is of the order $M_{D}/M_{R}\sim\sqrt{M_{\nu}/M_{R}}$, where $M_{R}=\sqrt{2}v_{R}{\bf f_{R}}$. Consequently, this mixing can vary significantly depending on the flavor structure of $\bf f_{R}$}. On the other hand, a small region of parameter space for $140\,\mathrm{MeV}\lesssim M_{N_{R}}\lesssim 500\,\mathrm{MeV}$ is excluded by recent analyses of meson decays [Alves:2023znq], particularly for $M_{W_{R}}=7\,\mathrm{TeV}$. However, this exclusion is unlikely to have a significant impact on our results. .
In Fig. 4, we also include the curves of $(T_{1/2}^{0\nu})^{-1}$ as a function of $|(f_{R})_{ee}|$ for $M_{W_{R}}=15\,\mathrm{TeV}$ in yellow. As we increase the value of $M_{W_{R}}$, the BSM effects are suppressed, $(T_{1/2}^{0\nu})^{-1}$ is smaller, and gets closer to the prediction in the standard mechanism. In particular, for $M_{W_{R}}=15\,\mathrm{TeV}$, the whole parameter region of $|(f_{R})_{ee}|$ satisfies the KamLAND-Zen constraint. Turning into the effects of the mixing angle, as shown in Eqs. (32), (35)–(37), $\zeta$ impacts the sub-amplitudes involving a right-handed neutrino exchange. In simple terms, compared with the $\zeta=0$ case, a non-zero mixing angle extends the parameter region where the light right-handed neutrino description is applicable and reduces the parameter region where the heavy right-handed neutrino description is relevant. By comparing the left and right panels in Fig. 4, we notice that the peaks of $(T_{1/2}^{0\nu})^{-1}$ are higher and further to the right for the cases of $\zeta\neq 0$, respectively, whereas the local minima are higher and closer to the peaks for the cases of $\zeta\neq 0$, respectively. We can now analyze the entire Fig. 3 in light of these findings. As expected, we noticed that the gap between the allowed regions narrows with increasing $W_{R}$ boson mass (from left to right), widens with non-zero $W_{L}-W_{R}$ mixing (from top to bottom), and disappears completely for $M_{W_{R}}=15\,\mathrm{TeV}$ for the current $0\nu\beta\beta$} decay constraints.
@@
Figure 6: Same as in Fig. 5, combined constraints from LNC and $0\nu\beta\beta$}-decay experiments on the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and the Yukawa coupling $|(f_{R})_{ee}|$ in the parity-violating LRSM in the normal hierarchy with $M_{W_{R}}=15\,\mathrm{TeV}$ for zero (left) and non-zero (right) $W_{L}-W_{R}$ mixing.
@@
To illustrate the complementary sensitivities of $0\nu\beta\beta$}-decay searches and LNC processes in both low-energy experiments and high-energy colliders, we present results based on current constraints and future prospects, as described in Sec. 3 and Sec. 4. The experimental sensitivities to $M_{\delta_{R}^{\pm\pm}}$ and $|(f_{R})_{ee}|$ are shown in Figs. 5 and 6 for zero and non-zero mixing angles. In each plot, the vertical axis on the right gives the values of $M_{N_{R}}$ that corresponds to $|(f_{R})_{ee}|$ using Eq. (44).
@@
In these two figures, the $0\nu\beta\beta$} decay exclusion limits are shown as red solid lines and correspond to the doubly-charged scalar (top diagonal), heavy right-handed neutrino (upper horizontal), and light right-handed neutrino (lower horizontal) contributions to the amplitude. They were obtained using the analytical expressions in Eqs. (45)–(47). The shaded red regions limited by those lines are excluded by the KamLAND-Zen experiment. Similar prospects for ton-scale experiments are included in dashed magenta lines. For comparison purposes, we can also calculate the $0\nu\beta\beta$} decay exclusion limit derived in Ref. [Dev:2018sel] with the left-right mixing angle $\zeta=0$. Rescaling their limit to accommodate the latest KamLAND-Zen results and modifying the corresponding values of $M_{W_{R}}$ [^7] Ref. [Dev:2018sel] assumed a fixed value of $v_{R}=5\sqrt{2}\,\mathrm{TeV}$. Under the assumption of $g_{R}=2g_{L}/3$, this corresponds to $M_{W_{R}}=2.2\,\mathrm{TeV}$, which has been excluded by the LHC searches at Run-2 phase [ATLAS:2023cjo, CMS:2022yjm, CMS:2021dzb]. one obtains $|(f_{R})_{ee}|/M_{\delta_{R}^{\pm\pm}}^{2}\lesssim 0.1\times(M_{W_{R}}/{\rm TeV})^{3}$, not visible in the range of Fig. 5. Notice how the use of an EFT description provides a stronger constraint in the $0\nu\beta\beta$}-decay sensitivity compared to the existing literature [Dev:2018sel].
@@
We also include collider searches in Figs. 5 and 6. The current indirect LEP exclusion limit for Bhabha scattering is shown as a solid orange line, with its corresponding exclusion region shaded in orange. The prospects for Bhabha scattering at the CEPC and the FCC-ee are shown as dashed dark green and light green lines, respectively, using the rescaling factor in Eq. (20). Notice that the difference in luminosities between the two accelerators makes the CEPC bound stronger. These collider searches will reach into the “funnel” portion of the parameter space that the $0\nu\beta\beta$} decay cannot access with current envisioned experiments. Direct searches for right-handed doubly-charged scalar $\delta_{R}^{\pm\pm}$ at hadron colliders are also shown for both current (LHC, grey solid line) and future (FCC-hh, grey dashed line) experiments, as discussed in Sec. 3. They are sensitive to $M_{\delta_{R}^{\pm\pm}}\lesssim 5\,\mathrm{TeV}$ irrespective of the Yukawa coupling $|(f_{R})_{ee}|$. If we conceive a scenario where a positive event in ton-scale $0\nu\beta\beta$} decay experiments is seen but both future leptons and hadrons colliders do not get a corresponding signal, then we can pin down the underlying $0\nu\beta\beta$}-decay mechanism to be dominated by the right-handed neutrino exchange and the corresponding parameter region would be a horizontal fringe of smaller $|(f_{R})_{ee}|$ in Figs. 5 and 6.
6 Conclusion